OpenStudy (anonymous):

someone help me on this problem 4y=5x-1 please 3x-2y=1

OpenStudy (anonymous):

it's a simple system of equations. I solve it so that I express x with the y in one equation or the other way around. 3x-2y=1 so -2y=1-3x, so 2y=3x-1 4y=5x-1 2*2y=5x-1 and then you replace 2y with 3x-1 2(3x-1)=5x-1 and you can do the rest :)

OpenStudy (anonymous):

who can help me on this question? -A characteristic of most “fad” diets include: A. inclusion of all food groups, utilizing a healthy eating plan C.inclusion of physical activity B.use of “magic” or “miracle” foods that burn fat D.diverse menu options